# Assembly Theory Reduced to Shannon entropy and Rendered Redundant by Naive Statistical Algorithms

### v7

@@ 2.2 Assembly Index is a Classical Compression Algorithm
The results in Fig. 3 accord with our mathematical proof demonstrating that the assembly index (and any variation based on their concept of ‘copy number’ based on counting identical copies in data) approximates, and is equivalent to, Shannon entropy via an $\mathbf{LZ}$ compression scheme and is, therefore, indistinguishable from Shannon entropy and $\mathbf{LZ}$ algorithms (of which LZW is one).
It follows that results previously reported by Assembly Theory are, therefore, given their relationship to and implementation of Shannon entropy via $\mathbf{LZ}$ compression, not unique. There is no fundamental or methodological advantage to renaming Shannon entropy as ‘assembly index’ or ‘Assembly Theory’, as we will show that the application of Shannon entropy and even LZW directly to the same data yields the same or better results.
2.3 ZBC Strings
@@
Their main argument in [1] is tested on Figs. 12 and 3. The contrast is evident between the range of values that Ai—and indeed LZW and Shannon entropy—produces for the fixed-length permutations of the ZBC base string, and that produced by BDM, a measure constructed on fundamentally different or extensive principles, as opposed to simply counting number of repetitions as Ai, LZW and Shannon entropy do–a fact which explains their similarities.
@@
Figure 2: Main figure: correlation plot between the most popular statistical compression algorithm LZW (behind ZIP and PNG) and the assembly index (Ai) for a growing ZBC sequence (length size of 100 characters) and its random permutations following the experiment in [1]. The colour code corresponds to the number of strings which produced the pair (Ai,LZW) during the experiments. The authors of Assembly Theory only reported the component with string size equal to 15 (highlighted in (b)), in their effort to ‘demonstrate’ the novelty of Ai compared to LZW [1] with a weak 0.25 correlation, without conducting or reporting on the full experiment. The apparent scattering of it is an artifact of the fixed scale. The quick asymptotic convergence to Spearman correlation 1 for multiple examples is provided in Table 1.
The analysis of Fig. 2 indicates that when the larger experiment is carried out, the full extent of the asymptotic convergence is revealed between Ai and LZW, with a correlation value of 1.00 to LZW. This is in complete agreement with the theoretical expectation of the mathematical proof of equivalence offered before, showing the equivalence of Ai with $\mathbf{LZ}$ and Shannon entropy [4]. Effectively, this means that any previous or future positive (or negative) results obtained using AT would be due to its equivalence to statistical compression and Shannon entropy. In order to arrive at the exact plot by using the equality with $\mathbf{LZ}$ and Shannon entropy, one has only to run LZW multiple times and pick the shortest path, or run Shannon entropy on multiple variable window lengths, but because this will require an heuristic as a result of the many ways in which it can be applied, just as for the Assembly index, the chances of producing the exact values and an identical plot data point by data point will be small. Yet, they will always be almost identical and fully convergent except for a vanishing number of outliers slightly deviating from the average. In other words, the difference may only be a small vanishing constant, given the heuristics chosen in the many implementations of the calculations of the Assembly index, heuristics which the authors have changed from paper to paper.
The claims in [1] wrongly suggest that the Assembly Index enjoys some sort of invariance theorem such as algorithmic (Kolmogorov) complexity (or $K$, guaranteeing convergence in values and independence of data representation [1]. This illustrates the multiple levels of misunderstanding and misleading statements in the papers on Assembly Theory. Although no invariance theorem holds for Ai, an invariance theorem guarantees convergence of $K_{T}$ values for different language representations $T$. When fixing the data representation, as the authors of AT do, neither $K$ nor any other measure depends on any change of data representation when applied to data that the authors of AT claim to be ‘empirical’ in nature, referring to chemical bond data (even as they have also suggested they can apply Ai to anything else, even text [20] which amounts simply to text compression with ZIP). Therefore, the arguments against $K$ or LZW in terms of data representation are not valid, as they have solved any potential problem by fixing the data representation.
@@
Figure 3: Density correlation plots for strings of growing length confirm the convergent asymptotic behaviour between the assembly index and LZW. Growing strings to the length of 100 with A and B showing random sequences, and the bottom plots C and D showing patterned sequences of repeated blocks. C shows a patterned block of ABCDE repeating, while D shows a ten-letter repeating block. The variance is lower in the random sequences, with growing size converging to very high correlations of 0.98-0.99, while in patterned sequences the correlations of the growing chain are high– about 0.9.
@@
The long-string asymptotic linear behaviour for Ai and LZW is shown in Fig. 4 indicating how the mean and standard deviation of the ratio of the logarithms of LZW and Ai behave and correlate.
@@
Figure 4: Asymptotic behaviour of (a) the mean and (b) the standard deviation of the ratio between log LZW and log Ai values for strings up to size 200 built from randomly picking characters from the ZBC string 10 000 times for each size.
@@
Table 1: Averaged Spearman correlations of growing patterned sequences (increasing copy number) between assembly index and other complexity measures
@@
Table 2: Convergence of averaged Spearman correlations of growing random sequences between assembly index (Ai) and various other complexity measures. Not surprisingly, because LZW, Entropy rate and BDM count exact copies in data as their most basic test for non-randomness, the convergence is fast and full. Ai is the baseline of these other algorithms. In addition, BDM also accounts for linear and algorithmic transformations (reversion, complementation, inversion, etc.) [8, 7].
@@
2.4 Clustering Molecules from Experimental Data
Fig. 5 on the right-hand side, shows four different clustering algorithms to visualise how well the mass spectra peak matrices (i.e., m/z ratio vs. a number of peaks) cluster by the three categories. As shown, hierarchical clustering, Gaussian Mixture Models (GMM), KMeans cluster, SVM, Random Forest (RF), KNN, and a neural network (FFNN) or multilayer perceptron, all but PCA separate the three categories.
Figure 5: Clustering and machine learning classification performance on mass spectrometry peak matrices from source data showing that any naive statistical or machine learning algorithm (for both transparent- and opaque-box approaches) can reproduce the results that Assembly Theory deems the ones unifying biology and physics [3].
In Fig. 5, from A to D on the left-hand side, we see the ROC curves of four ML algorithms classifying the organic and inorganic molecules (N=30) from the Marshall et al. [5] paper’s Fig. 4 mass spectrometry mixtures into 3 categories: Dead, Inorganic, and Biological. SVM had the highest area under the curve (AUC), with 0.75, and a classification accuracy of 0.86. SVM suggests linear hyperplane separability.
The ML predictions are further supported by the ROC curves in Figure 6. Fig. 6 demonstrates that 1D-BDM and LZW compression measures significantly outperform MA as complexity measures in distinguishing between inorganic, dead, and biological molecules from [5] mass spectrometry data. Specifically, the AUC scores for BDM and LZW approaches are perfect or near-perfect (AUC = 1.00), while MA struggles with considerably lower AUC scores: within the 0.43-0.77 range with ML classification. This underperformance suggests that MA is suboptimal for detecting and quantifying biosignatures, possibly due to its less robust representation of molecular complexity in comparison to complexity measures from the AIT framework, like BDM and LZW, which more effectively capture the intricate patterns needed to differentiate between these categories. Consequently, MA (Ai) may not be well-suited for the study’s objective of identifying unique biosignatures of life based on molecular complexity and its multi-scale organisation.
Figure 6: ROC curves comparing mass spectrometry data of organic and inorganic molecules by way of comparing the measures of 1D-BDM and MA. LZW, and other compression schemes like RLE and Huffman, had similar values close to the evaluated ML performance of BDM (AUC scores of 0.89 for MLP and 1.00 for KNN and RF). It is clearly shown that BDM (and, similarly, other AIT measures) outperforms MA (the values attributed to the environmental samples’ mass spectrometry data in Marshall et al. [5] Figure 4) using basic machine learning algorithms. Likewise, LZW compression also outperforms MA, with a perfect AUC (1.00) using FFNN and RF and a better AUC than MA for KNN (0.69). MA has poor AUC scores of 0.47 for MLP and 0.43 in KNN and RF-based classification. All this even gives AT and MA the advantage of using their own small choice of molecules [5] rather than the exhaustive experiment we reported years before [12] on more than 15 000 compounds.
2.5 Strong expected correlation of Molecular Assembly Index with LZW
To explain how almost any naive algorithm can reproduce the results reported by the authors of Assembly Theory in [5, 3], we took the data from [5] and reproduced their analysis in Fig. 7 by exploiting the linear relationships found between LZW, the length of the InChI strings for molecules and MS2 data, respectively.
We say that the length of the InChI string is the number of characters existing after the initial prefix “InChI=1S/”. We also plot in Fig. 7 the relation to MA (that is, Ai), as presented by Marshall et al. [5] in their Fig.3.
Figure 7: Quantile regressions considering 5%, 50 % and 95% quantiles and Pearson correlation (corr) between MA (top), Length of InChI string (middle) and LZW of InChI string (bottom) and MS2 using the data from [5]. The shaded area represents a 90 % quantile interval. For any of the metrics, linear relationships with similar prediction capabilities are observed. This directly undercuts the AT authors’ claims that one of the main advantages of MA (or, equivalently, Ai) is that it could be experimentally measured, since all the other metrics (LZW and the Length of the InChI codes) can also be estimated from experimental MS2 data directly, as we have done here. In contradiction to what the authors claim in [1], the authors of AT have failed at controlling for molecular length, which drives their index. The authors of AT have thus far failed to show a single example in which $\mathbf{LZ}$ or Shannon entropy cannot reproduce any results produced by Ai or AT, literally meaning Ai and AT are $\mathbf{LZ}$ or Shannon entropy with extra steps.
One might ask why plotting the relation between MA and the length of InchI strings is of interest. This is a consequence of the asymptotic behaviour observed for compression rates in $\mathbf{LZ}$ schemes, which tend to a constant for large strings. Since Ai belongs to the $\mathbf{LZ}$ family [4], Ai should be linearly related to the length of the base string for which it is calculated. Thus, it would be natural to observe some type of linear relation, which was verified in Fig. 8.
Figure 8: Linear relation between MA and the Length of the InChI strings of molecules using the data from [5]. This was conducted on the same small set of molecular compounds picked by the authors [5], unlike the exhaustive experiment we reported years before [12]. The authors of AT did not control for molecular length, which would have deleted any effect of Ai as it does not pick up any other signal.
Figure 9: Values of several complexity measures for Small Molecules and Peptides, using the data from [5]. This is the calibration experiment performed in [5] before moving to classify organic vs nonorganic compounds. While in [12] we normalised by molecular length and compared the efficiency of the organic vs inorganic separation on over 15 000 compounds reported (the full ChemicalData[] database from the Wolfram language), not a small selection [5] of only about 130 compounds, with no comparisons to any other complexity measure that would have disclosed similarity or superiority at separating the classes. MA is one of the many versions of Ai across the authors’ papers for application to molecules, implementing their ‘copy-number’ hypothesis.
From our numerical experiments it is clear that LZW and Ai produce results that cannot be told apart. In Fig. 3, a clear linear asymptotic tendency is observed between LZW and Ai values. This suggests that one can be easily obtained from the other by simple algebraic operations that involve considering certain dictionary elements, basically a small constant term. The plots in Fig. 3 demonstrate that the variance of Ai is much greater than that of LZW. This comes from the rigid structure that arises from the LZW algorithm (always move from left to right, in a single pass) but many other $\mathbf{LZ}$ algorithms implement variations, of which Ai is one.
Fig. 4 shows that the mean value of (log LZW/log Ai) tends to a constant and that the standard deviation of (log LZW/log Ai) tends to 0 for large strings. This indicates that the random variable (log LZW/log Ai) tends to a single-bar histogram, indicating that this ratio tends to a constant without variability. This is an empirical demonstration that the two values converge in alignment with the mathematical proof of equivalence we provided in [4].
Figs. 7 demonstrate that the claimed link that Ai has with experimental results is also observed for LZW and even for the length of the InChI strings of the molecules analysed. Thus, if we follow the same rationale as Marshall et al.[5], it could be said that LZW of the InChI strings, and even on their lengths only, can be experimentally measured by molecules’ mass spectra. This lacks, of course, a physical/chemical explanation, which is never provided by AT, beyond circular arguments [4].
Also, by taking into consideration the results from the experiments in Fig. 3, one could investigate how the Ai values are related to the length of the InChI strings. Fig. 8 displays a remarkably strong linear relationship, with a $R^{2}$ of over 0.9 and Pearson correlation of 0.95. This indicates that Ai is a proxy for the length of InChI strings, revealing that these strings already present a great share of the “concatenation” information for the molecules.
Regarding the threshold suggested by [5], that according to the authors separates molecules produced by non-living systems from molecules produced by living ones, it is claimed that non-living systems do not produce molecules with MAs (Ai) higher than 15. This has already been found to be erroneous by two other independent groups  [10, 11]. However, in their defence [1] in this regard, the authors suggest that one has to filter out those examples that prove the Ai hypothesis wrong because they are too artificial, while at the same time, the same authors claim Ai is completely agnostic and capable of detecting any form of life in the universe [5, 3].
This claim follows from [5, Fig. $4$], where it is asserted that only molecules produced by living organisms (and some “Blinded” samples from NASA/other sources that were not made fully available in their paper’s dataset, bypassing Nature’s rules and review procedures) have MAs (Ai) higher than 15. By using their own data, we reproduced their [5, Fig. $4$] in our Fig. 10, taking the results for Ai at face value as they were not reproducible, given that despite their claims to the contrary, they did not include all the data. The threshold has been already challenged and found to be in error twice by other independent groups [10, 11], but here we show clearly that their assembly index (Ai) is not picking up any special signal from life.
The first fact worth noting is that Marshall et al. [5] did not use any of the quantile regressions they indicated in their [5, Fig. $3$] to estimate MA (Ai) values from MS2 peaks (i.e., they did not do any of the 5%, 50 % and 95% quantiles, which are also reproduced in Fig. 7). By experimenting with the quantile value, it was possible to determine that they used a 14% quantile regression on the data to build their linear relationship. Then, by using this linear relationship and the measure MS2 peak data, Fig. 10 was obtained. This choice was not deeply explored in their paper, which indicates a serious methodological gap.
If we follow the same rationale as in [5], we could say that molecules produced by non-living systems could not have estimated InChI string lengths greater than 100 characters. This illustrates why their claim is neither methodologically robust nor correct. Thus, the analysis of Figs. 7 and 10 shows that the organic vs inorganic separation claimed in [5] is driven by molecular length and not any other feature captured by their (molecular) assembly index or their copy number, beyond triviality.
We reported all this years before AT in [12], using a much better conducted experiment involving over 15,000 compounds, comparing to several measures and several data representations, and without offering the results as proof that we could unify physics and biology [3]. The rationale behind Assembly Theory is equivalent to finding any variable linearly correlated to MS2 and then using an arbitrary threshold to separate the molecules, which is without any physical/chemical foundation and does not withstand basic scrutiny, despite the characterisation of the results as extraordinary and the claim to have found a new ‘Theory of Everything’ [3]. In fact, this is exactly the opposite of what they claim to have done. They have unwittingly fallen into the classic correlation as causation trap, where one finds a correlated variable and claims a causal effect.
For example, this is evinced in later attempts to either bypass or address those criticisms, as in [24]. Despite the decades of research showing the application of information theory and compression methods to applied sciences, assembly theory’s proponents keep offering up the fallacy— as in the abstract of [24]— that compression methods are not efficiently applicable, or in general that algorithmic-complexity-based measures cannot be applied to empirical investigation. In [24], the Molecular Assembly Index (based on the assembly index) was shown to correlate (with a Pearson’s of $0.88$ for their best result) to an ensemble of experimentally measured data. If this correlation “demonstrates” that Assembly Theory is an experimental measure, then it follows that the strong correlation we have shown would also “demonstrate” that one can measure assembly index by using LZW. Contrary to implying equivalences from naive correlation arguments, in [4] we have formally demonstrated that the assembly index method is a (statistical) compression method, empirically corroborated in [2] and in the present article.
Since the relation between MS2 peaks and MA (Ai) is linear (the same for InChI lengths), what was actually proposed in one of the many papers in question [5] is that molecules produced by living and non-living systems can simply be distinguished by the number of their MS2 peaks (which were measured and not estimated). Fig. 10 also presents the plot of the raw data from [5], where it is easily seen that their threshold of 15 for estimated MA values is actually a proxy of the threshold of about 25 for directly measured MS2 peak data. This does not seem to be a physically/chemically acceptable claim, as discussed further in the Supplementary Information.

### v10

@@ 2.2.1 Analysis on Synthetic Strings
The results in this section accord with our mathematical proof demonstrating that the assembly index (and any variation based on their concept of ‘copy number’ based on counting identical copies in data) approximates, and is equivalent to, Shannon entropy via an LZ compression scheme and is, therefore, indistinguishable from Shannon entropy and LZ algorithms (of which LZW is one). It follows that results previously reported by Assembly Theory are, therefore, given their relationship to and implementation of Shannon entropy via LZ compression, not unique.
@@
The authors of AT offer a sequence of very simple examples as evidence in their favour. The first consists of the letters Z, B, and C arranged in long repetitive or random strings. Their main argument [4] is tested in Figs. 1 and 2. The contrast is evident between the range of values that Ai—and indeed LZW and Shannon entropy—produces for the fixed-length permutations of the ZBC base string, and that produced by BDM, a measure constructed on fundamentally different or extensive principles, as opposed to simply counting number of repetitions as Ai, LZW and Shannon entropy do—a fact which explains their similarities.
@@
Figure 1: Main figure: correlation plot between the most popular statistical compression algorithm LZW (behind ZIP and PNG) and the assembly index (Ai) for a growing ZBC sequence (length size of 100 characters) and its random permutations following the experiment in [4]. The colour code corresponds to the number of strings which produced the pair (Ai,LZW) during the experiments. The authors of Assembly Theory only reported the component with string size equal to 15 (highlighted in (b)), in their effort to demonstrate the novelty of Ai compared to LZW [4] with a weak 0.25 correlation, without conducting or reporting on the full experiment. The apparent scattering of it is an artifact of the fixed scale. The quick asymptotic convergence to Spearman correlation 1 for multiple examples is provided in Table 1 and its undisputed mathematical equivalence in [17]. Unlike the claim in [4], the assembly index correlation does not weaken with object size 1.
The analysis of Fig. 1 indicates that when the experiment is carried out only slightly further, the full picture and extent of the asymptotic convergence between Ai and LZW is revealed. This is in complete agreement with the theoretical expectation of the mathematical proof we offered before establishing the equivalence of Ai with Shannon entropy [17] (via LZ). Effectively, this means that any previous or future positive (or negative) results obtained using AT would be due to its equivalence to statistical compression and Shannon entropy and not because of any particularity or advantage of AT or Ai. Due to the heuristics introduced in the calculation of Ai by the authors of AT, some deviations from the diagonal can be observed with the vanishing number of outliers as a function of length. Those variations do not prove the difference beyond local anomalies and do not grant AT or Ai any of the features the authors have represented it with, including claims related to the abilities of explaining selection and evolution.
@@
Figure 2: Density correlation plots for strings of growing length confirm the convergent asymptotic behaviour between the assembly index and LZW. A and B are random sequences of length up to 100 while C and D are patterned sequences of repeated blocks of up to 100 characters. C shows a patterned block of ABCDE letters repeating, while D shows a ten-letter repeating block. The variance is lower in the random sequences, with growing size converging to very high correlations of 0.98-0.99, while in patterned sequences the correlations of the growing chain are high at about 0.9. Even when visually it may appear that the scattering increases for patterned strings, this is a visual artifact from the number of greater block lengths. As shown in the numerical Table 1, the correlation in all cases converges quick to 1 as a function of length as shown in Figs. 3(a),3(b). But the exercise was futile given that we had already proven the mathematical equivalence in [17]. Unlike the claim in [4], the assembly index correlation does not weaken with object size as shown in Table 1.
@@
The linear asymptotic behaviour of the long string for Ai and LZW is shown in Fig. 3 indicating how the mean and standard deviation of the ratio of the logarithms of LZW and Ai behave and correlate. In [17], however, we had already proven the mathematical equivalence the authors did not address.
@@
Figure 3: Asymptotic behaviour of (a) the mean and (b) the standard deviation of the ratio between log LZW and log Ai values for strings up to size 200 built from randomly picking characters from the ZBC string 10,000 times for each size. These plots quantify the Assembly index (Ai) full convergence to LZW.
@@
Table 1: Averaged Spearman correlations of growing patterned sequences (increasing copy number) between assembly index and other complexity metrics that the original authors should have compared with under good scientific practice but have always avoided to. Correlation is true for short strings, even with under 15 copies of the same element (the threshold for life, according to AT). The Block Decomposition Method (BDM), was introduced almost a decade before AT. Unlike the claim in [4], the assembly index correlation does not weaken with object size because AT with Ai instantiates Shannon entropy through LZ. However, BDM has been proven to go beyond Shannon entropy and therefore beyond Ai. However, it takes as a baseline what Ai and Shannon entropy do [17] which is the ‘counting repetitions’ baseline that is the most trivial and considered by virtually every other complexity measure [6].
@@
Table 2: Convergence of averaged Spearman correlations of growing random sequences between assembly index (Ai) and various other complexity measures. Not surprisingly, because LZW, Entropy rate and BDM count exact copies in data as their most basic test for nonrandomness, the convergence is fast and full. Ai is the baseline of these other algorithms. In addition, BDM also accounts for linear and algorithmic transformations (reversion, complementation, inversion, etc.) [7, 6].